OpenStudy (precal):

The number of gallons of water in a swimming pool t minutes after the pool has started to drain is Q(t) = 50(20 - x)2. How fast is the water running out at the end of 12 minutes?

OpenStudy (precal):

I got -800 gal/min but the key states +800 gal/min What am I doing wrong?

OpenStudy (precal):

calculus problem

OpenStudy (phi):

you found the rate as -800 g/min in English you could say this means the water leaves at a rate of 800 g/min

OpenStudy (dumbcow):

ok so you have to find Q'(12)...... the derivative function where t=12 use chain rule....inside = 20-t , derivative of inside is -1 outside = 50u^2 , derivative = 100u \[Q'(t) = (-1)100(20-t)\] plug in t=12 \[Q'(12) = -100(20-12) = -800\]

OpenStudy (dumbcow):

oh crap i didnt need to do that :{

OpenStudy (precal):

ok so I am correct the negative represent the water leaving just like in related rates when we add negative and positive signs when needed

OpenStudy (precal):

thanks for helping me

OpenStudy (phi):

yes, your math is correct (as confirmed by dumb cow). So it comes down to semantics water leaving versus water coming in at a negative rate.